# EdiText: Controllable Coarse-to-Fine Text Editing with Diffusion Language Models

### v1

@@ Abstract
We propose EdiText, a controllable text editing method that modify the reference text to desired attributes at various scales. We integrate an SDEdit-based editing technique that allows for broad adjustments in the degree of text editing. Additionally, we introduce a novel fine-level editing method based on self-conditioning, which allows subtle control of reference text. While being capable of editing on its own, this fine-grained method, integrated with the SDEdit approach, enables EdiText to make precise adjustments within the desired range. EdiText demonstrates its controllability to robustly adjust reference text at broad range of levels across various tasks, including toxicity control and sentiment control.
@@ 1 Introduction
Editing methods have also emerged in the text domain Du et al. (2022); Mallinson et al. (2020); Malmi et al. (2019); Pryzant et al. (2020). Although autoregressive (AR) language models Brown et al. (2020); Raffel et al. (2020); Touvron et al. (2023) generally outperform non-autoregressive (NAR) counterparts Clark et al. (2020); Devlin et al. (2019); Li et al. (2022); Liu et al. (2019), NAR models have shown strengths in controlled generation Li et al. (2022); Mireshghallah et al. (2022), where shifting to the target distribution can be easily achieved with relatively small amount of data and training. The ability of NAR models to facilitate such shifts can also be effectively applied to text editing. For example, one might need an edit to shift the overall tone of a report from informal to formal, or perform a local edit to change only specific words. In line with these possibilities, a number of studies Horvitz et al. (2024); Mallinson et al. (2020); Mireshghallah et al. (2022); Pryzant et al. (2020) have adapted NAR-based controlled generation methods to text editing. Although these models are capable of reflecting the target attributes in text to some extent, their editing techniques either lack the ability to control the degree of editing or are limited to a narrow range, which restricts their versatility.
@@
Recently, the advancement of diffusion models has extended into the discrete text domain, which can be categorized into two main types Zou et al. (2023): discrete diffusion Austin et al. (2021); He et al. (2023); Hoogeboom et al. (2021) and embedding diffusion Gao et al. (2022); Gong et al. (2023); Li et al. (2022). Embedding diffusion models, specifically, map discrete tokens to continuous embeddings and model them through a Gaussian diffusion process, able to leverage various advancements from the diffusion models in other continuous domains Li et al. (2022); Lovelace et al. (2023); Zhang et al. (2024a). While this advantage demonstrates the potential of applying several techniques from continuous diffusion models Chen et al. (2023); Dhariwal and Nichol (2021); Ho and Salimans (2021), diffusion-based text editing has not yet been sufficiently explored.
@@ 2 Related Works
Diffusion Language Model. Diffusion models have achieved remarkable performance in vision and audio tasks—especially in generation and editing. Building on these successes, there has been a growing interest in extending the diffusion framework to the text domain. A number of studies Chen et al. (2023); Lovelace et al. (2023); Hoogeboom et al. (2021); Austin et al. (2021); Sahoo et al. (2024); Lou et al. (2024) have adopted diffusion framework for text generation and confirmed their effectiveness. Moreover, several works Li et al. (2022); Zhang et al. (2024a); Gong et al. (2023) have highlighted the controllability of these models in constrained scenarios-such as tasks involving parse trees or fixed sequence lengths-by leveraging various guidance techniques Ho and Salimans (2021); Dhariwal and Nichol (2021) tailored for diffusion frameworks. Despite these strengths, the potential of diffusion language models for text editing—a task that could benefit significantly from their broad controllability—remains underexplored.
Controlled Text Generation. Many studies have focused on methods to control the generation process to align with user intentions. One prominent task, controlled text generation, seeks to produce text that adheres to specific constraint (i.e. fixed sequence length, parse-tree, or specific target attributes). While some approaches leverage AR models for controlled text generation Carlsson et al. (2022); Yang et al. (2023), NAR models Li et al. (2022); Zhang et al. (2024b); Mireshghallah et al. (2022) often require fewer parameters and benefit from a parallel sampling process. This parallelism enables NAR models to consider the entire context of the textual input, facilitating smoother distribution shifts during the sampling process.
Text Editing. A task similar to controlled text generation is text editing, which involves modifying a given reference text to align with the user’s intent. Text editing differs from controlled text generation in that it focuses on transforming existing content based on specific goals, rather than generating entirely new text from scratch. Several studies have proposed reference-based editing methods Malmi et al. (2019); Du et al. (2022); Laban et al. (2023); Raheja et al. (2023), while some are also known as controlled paraphrase generation Ogasa et al. (2024); Dehghani et al. (2021); Yang et al. (2022) which focuses on specific constraints. However, these approaches generally lack mechanisms to control the degree of editing, making it difficult to achieve the desired level of transformation.
@@ 3 EdiText
In this section, we introduce EdiText, a text editing framework which utilizes a diffusion model capable of performing both fine-grained and coarse-grained editing. We adopt Latent Diffusion for Language Generation (LD4LG) Lovelace et al. (2023), an embedding diffusion model, as the backbone of our framework to support text editing where the discrete text is transformed into continuous data, allowing diffusion-based editing methods, as described in Section 3.1. Inspired by the image editing technique SDEdit Meng et al. (2022), we present a method for controlling text editing at a coarse level in Section 3.2. Furthermore, to support fine-grained control of the extent of editing, we propose a novel self-conditioning-based approach Chen et al. (2023) in Section 3.3. By combining these elements, EdiText is capable of editing text of varying attributes at different levels of granularity.
@@ 3.1 Latent Diffusion for Language Generation
To model the distribution of the latent representation, Lovelace et al. (2023) first define a forward diffusion process which gradually transforms $x_{0}=x$ into random noise $x_{T}\sim N(0,I)$ over timesteps $T$. This forward process is defined by a noise schedule $\alpha_{t}$, and is utilized to compute the corrupted representation $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t}$ at an intermediate timestep $t\in[1,T]$, where $\epsilon_{t}$ is a noise vector that follows a standard normal distribution. To sample the latent representation $x_{0}$, Lovelace et al. (2023) follow the reverse trajectory of the pre-defined forward process learned by a diffusion model. They train a model $x_{\theta}$ to predict $x_{0}$ from noisy data $x_{t}$, which is utilized to generate the latent representation $x_{0}$ from noise $x_{T}$. The equations used for training and sampling procedure of LD4LG, as described above, are as follows:
$L(\theta)$ $={\mathbb{E}_{t,x_{0},\epsilon_{t}}[\lambda_{t}^{-1}\lVert x_{\theta}(x_{t},t)-x_{0}\rVert_{2}^{2}]}],$ (1)
@@
$\epsilon_{\theta}(x_{t},t)$ $=\lambda_{t}^{-0.5}(x_{t}-\sqrt{\alpha_{t}}x_{\theta}(x_{t},t))$ (3)
where $\epsilon_{\theta}(x_{t},t)$ is the reparameterized version of $x_{\theta}(x_{t},t)$ and $\lambda_{t}=1-\alpha_{t}$.
@@
To improve sample quality, LD4LG leverages the self-conditioning Chen et al. (2023), which uses the previous timestep’s prediction as an additional condition for the diffusion model at current timestep. Specifically, during training at timestep $t$, the diffusion model $x_{\theta}(x_{t},t,\emptyset)$ is trained using Eq. 1 with a pre-defined probability $p$. Here, $\emptyset$ indicates that, as in a typical diffusion model, no extra condition is provided beyond the current timestep $t$ and the corrupted data $x_{t}$ when predicting $x_{0}$. Conversely, with probability $1-p$, the model employs the prediction $\hat{x}_{0,t}=x_{\theta}(x_{t},t,\emptyset)$ as a condition, thereby training $x_{\theta}(x_{t},t,\hat{x}_{0,t})$ to predict $x_{0}$. During self-conditioning-based training, gradients do not propagate back through the condition $\hat{x}_{0,t}$, and the loss function is defined as follows:
$L={\mathbb{E}_{t,x_{0},\epsilon_{t}}[(1-\alpha_{t})^{-1}\lVert x_{\theta}(x_{t},t,\hat{x}_{0,t}))-x_{0}\rVert_{2}^{2}]}].$ (4)
To better adapt LD4LG for text editing, we train a class-conditional LD4LG that accepts a desired attribute as condition. Given a class label $c$, we replace $x_{\theta}(x_{t},t,\emptyset)$ and $x_{\theta}(x_{t},t,\hat{x}_{0,t})$ with $x_{\theta}(x_{t},t,\emptyset,c)$ and $x_{\theta}(x_{t},t,\hat{x}_{0,t},c)$ during training. This change lets LD4LG generate text with the target condition, thereby extending its application to text editing by incorporating our proposed editing methods, as described in the following section.
3.2 SDEdit-based Coarse-Level Editing
@@
Inspired by these aspects demonstrated in the vision domain, we adopt SDEdit into the text domain, enabling the editing of textual data with variations ranging from very slight changes to substantial changes that diverge significantly from the reference text by controlling $t_{CE}$. This allows for a wide range of text manipulation, facilitating coarse-level text editing. Our SDEdit-based coarse-level editing method is described in Algorithm 1.
3.3 Self-Conditioning-based Fine-Level Editing
@@
Unlike SDEdit, which starts denoising at intermediate timesteps $t_{CE}$ to provide the reference text information more directly, the self-conditioning-based method indirectly feeds the reference text as a condition during the initial sampling steps, resulting in smaller variations in the level of editing. Although SDEdit provides coarse-level editing and self-conditioning offers fine-level adjustments, EdiText can combine both approaches to enhance controllability. It first sets the overall extent of editing and then applies fine-grained modifications to achieve the desired effect. Details on the integrated EdiText algorithm are in the Appendix A.1.
@@ 4.1 Datasets
4.2 Implementation Details
4.3 Models
4.4 Evaluation Metrics
5 Results

### v2

@@ Abstract
We propose EdiText, a controllable text editing method that modifies the reference text to desired attributes at various scales. We integrate an SDEdit-based editing technique that allows for broad adjustments in the degree of text editing. Additionally, we introduce a novel fine-level editing method based on self-conditioning, which allows subtle control of reference text. While being capable of editing on its own, this fine-grained method, integrated with the SDEdit approach, enables EdiText to make precise adjustments within the desired range. EdiText demonstrates its controllability to robustly adjust reference text at a broad range of levels across various tasks, including toxicity control and sentiment control.
@@ 1 Introduction
Editing methods have also emerged in the text domain Du et al. (2022); Mallinson et al. (2020); Malmi et al. (2019); Pryzant et al. (2020). Although autoregressive (AR) language models Brown et al. (2020); Raffel et al. (2020); Touvron et al. (2023) generally outperform non-autoregressive (NAR) counterparts Clark et al. (2020); Devlin et al. (2019); Li et al. (2022); Liu et al. (2019), NAR models have shown strengths in controlled generation Li et al. (2022); Mireshghallah et al. (2022), where shifting to the target distribution can be easily achieved with a relatively small amount of data and training. The ability of NAR models to facilitate such shifts can also be effectively applied to text editing. For example, one might need an edit to shift the overall tone of a report from informal to formal, or perform a local edit to change only specific words. In line with these possibilities, a number of studies Horvitz et al. (2024); Mallinson et al. (2020); Mireshghallah et al. (2022); Pryzant et al. (2020) have adapted NAR-based controlled generation methods to text editing. Although these models are capable of reflecting the target attributes in text to some extent, their editing techniques either lack the ability to control the degree of editing or are limited to a narrow range, which restricts their versatility.
@@
Recently, the advancement of diffusion models has extended into the discrete text domain, which can be categorized into two main types Zou et al. (2023): discrete diffusion Austin et al. (2021); He et al. (2023); Hoogeboom et al. (2021); Sahoo et al. (2024); Shi et al. (2024) and embedding diffusion Gao et al. (2022); Gong et al. (2023); Li et al. (2022). Embedding diffusion models, specifically, map discrete tokens to continuous embeddings and model them through a Gaussian diffusion process, able to leverage various advancements from the diffusion models in other continuous domains Li et al. (2022); Lovelace et al. (2023); Zhang et al. (2024a). While this advantage demonstrates the potential of applying several techniques from continuous diffusion models Chen et al. (2023); Dhariwal and Nichol (2021); Ho and Salimans (2021), diffusion-based text editing has not yet been sufficiently explored.
@@ 2 Related Works
Diffusion Language Model. Diffusion models have achieved remarkable performance in vision and audio tasks—especially in generation and editing. Building on these successes, there has been a growing interest in extending the diffusion framework to the text domain. A number of studies Austin et al. (2021); Chen et al. (2023); Hoogeboom et al. (2021); Lou et al. (2024); Lovelace et al. (2023); Sahoo et al. (2024); Shi et al. (2024) have adopted diffusion framework for text generation and confirmed their effectiveness. Moreover, several works Gong et al. (2023); Li et al. (2022); Zhang et al. (2024a) have highlighted the controllability of these models in constrained scenarios-such as tasks involving parse trees or fixed sequence lengths-by leveraging various guidance techniques Dhariwal and Nichol (2021); Ho and Salimans (2021) tailored for diffusion frameworks. Despite these strengths, the potential of diffusion language models for text editing—a task that could benefit significantly from their broad controllability—remains underexplored.
Controlled Text Generation. Many studies have focused on methods to control the generation process to align with user intentions. One prominent task, controlled text generation, seeks to produce text that adheres to specific constraints (i.e. fixed sequence length, parse-tree, or specific target attributes). While some approaches leverage AR models for controlled text generation Carlsson et al. (2022); Yang et al. (2023), NAR models Li et al. (2022); Mireshghallah et al. (2022); Zhang et al. (2024b) often require fewer parameters and benefit from a parallel sampling process. This parallelism enables NAR models to consider the entire context of the textual input, facilitating smoother distribution shifts during the sampling process.
Text Editing. A task similar to controlled text generation is text editing, which involves modifying a given reference text to align with the user’s intent. Text editing differs from controlled text generation in that it focuses on transforming existing content based on specific goals, rather than generating entirely new text from scratch. Several studies have proposed reference-based editing methods Du et al. (2022); Malmi et al. (2019); Laban et al. (2023); Raheja et al. (2023), while some are also known as controlled paraphrase generation Dehghani et al. (2021); Ogasa et al. (2024); Yang et al. (2022) which focuses on specific constraints. However, these approaches generally lack mechanisms to control the degree of editing, making it difficult to achieve the desired level of transformation.
Figure 2: Overview of the proposed EdiText framework. (a) We employ an embedding diffusion model as the backbone language model for EdiText (Section 3.1). (b) The two editing methods with different degree of editing—Coarse-Level Editing and Fine-Level Editing—are described in Sections 3.2 and 3.3, respectively.
@@ 3 EdiText
In this section, we introduce EdiText, a text editing framework which utilizes a diffusion model capable of performing both fine-grained and coarse-grained editing. We adopt Latent Diffusion for Language Generation (LD4LG) Lovelace et al. (2023), an embedding diffusion model, as the backbone of our framework to support text editing where the discrete text is transformed into continuous data, allowing diffusion-based editing methods, as described in Section 3.1. Inspired by the image editing technique SDEdit Meng et al. (2022), we present a method for controlling text editing at a coarse level in Section 3.2. Furthermore, to support fine-grained control of the extent of editing, we propose a novel self-conditioning-based approach Chen et al. (2023) in Section 3.3. By combining these elements, EdiText is capable of editing text of varying attributes at different levels of granularity. The overview of our proposed editing framework, EdiText, is illustrated in Figure 2.
@@ 3.1 Latent Diffusion for Language Generation
To model the distribution of the latent representation, Lovelace et al. (2023) first define a forward diffusion process which gradually transforms $x_{0}=x$ into random noise $x_{T}\sim N(0,I)$ over timesteps $T$. This forward process is defined by a noise schedule $\alpha_{t}$, and is utilized to compute the corrupted representation $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t}$ at an intermediate timestep $t\in[1,T]$, where $\epsilon_{t}$ is a noise vector that follows a standard normal distribution. To sample the latent representation $x_{0}$, Lovelace et al. (2023) follow the reverse trajectory of the pre-defined forward process learned by a diffusion model. Unlike conventional diffusion approaches that model the noise vector $\epsilon_{t}$ and reconstruct data from noisy samples, they employ a reparametrization strategy: they train a network $x_{\theta}$ to predict the clean latent $x_{0}$ directly from noisy input $x_{t}$.
The training objective for LD4LG minimizes the following loss:
$L(\theta)$ $={\mathbb{E}_{t,x_{0},\epsilon_{t}}[\lambda_{t}^{-1}\lVert x_{\theta}(x_{t},t)-x_{0}\rVert_{2}^{2}]}.$ (1)
where $\lambda_{t}=1-\alpha_{t}$.
At inference time, LD4LG iteratively denoise according to
@@
$\epsilon_{\theta}(x_{t},t)$ $=\lambda_{t}^{-0.5}(x_{t}-\sqrt{\alpha_{t}}x_{\theta}(x_{t},t)).$ (3)
where $\epsilon_{\theta}(x_{t},t)$ is the reparameterized version of $x_{\theta}(x_{t},t)$.
@@
To improve sample quality, LD4LG leverages the self-conditioning Chen et al. (2023), which uses the previous timestep’s prediction as an additional condition for the diffusion model at current timestep. During training, the diffusion model alternates between two distinct modes with probability $p$. Specifically, at timestep $t$, the diffusion model $x_{\theta}(x_{t},t,\emptyset)$ is trained using Eq. 1 with a pre-defined probability $p$. Here, $\emptyset$ indicates that, as in a typical diffusion model, no extra condition is provided beyond the current timestep $t$ and the corrupted data $x_{t}$ when predicting $x_{0}$. Conversely, with probability $1-p$, the model employs the prediction $\hat{x}_{0,t}=x_{\theta}(x_{t},t,\emptyset)$ as a condition, thereby training $x_{\theta}(x_{t},t,\hat{x}_{0,t})$ to predict $x_{0}$. During self-conditioning-based training, gradients do not propagate back through the condition $\hat{x}_{0,t}$, and the loss function is defined as follows:
$L={\mathbb{E}_{t,x_{0},\epsilon_{t}}[(1-\alpha_{t})^{-1}\lVert x_{\theta}(x_{t},t,\hat{x}_{0,t})-x_{0}\rVert_{2}^{2}]}.$ (4)
@@
To better adapt LD4LG for text editing, we train a class-conditional LD4LG that accepts a desired attribute as condition. Given a pre-defined class label $c$, i.e. toxifying or detoxifying label in a toxicity task, we replace $x_{\theta}(x_{t},t,\emptyset)$ and $x_{\theta}(x_{t},t,\hat{x}_{0,t})$ with $x_{\theta}(x_{t},t,\emptyset,c)$ and $x_{\theta}(x_{t},t,\hat{x}_{0,t},c)$ during training. This change lets LD4LG generate text with the target condition, thereby extending its application to text editing by incorporating our proposed editing methods, as described in the following section.
3.2 SDEdit-Based Coarse-Level Editing
@@
Inspired by these aspects demonstrated in the vision domain, we adopt SDEdit into the text domain, enabling the editing of textual data with variations ranging from very slight changes to substantial changes that diverge significantly from the reference text by controlling $t_{CE}$. This allows for diverse text manipulation, as reflected in wider retention and reflection scores, facilitating coarse-level text editing. Our SDEdit-based coarse-level editing method is described in Algorithm 1.
3.3 Self-Conditioning-Based Fine-Level Editing
Unlike SDEdit, which starts denoising at intermediate timesteps $t_{CE}$ to provide the reference text information more directly, the self-conditioning-based method indirectly feeds the reference text as a condition during the initial sampling steps, resulting in smaller variations in the level of editing. Although SDEdit provides coarse-level editing and self-conditioning offers fine-level adjustments, EdiText can combine both approaches to enhance controllability. It first sets the overall extent of editing and then applies fine-grained modifications to achieve the desired effect. Further details on both the coarse-level and fine-level algorithms, as well as the integrated algorithm, are in Appendix A.1.
@@ 4.1 Datasets
4.2 Implementation Details
4.3 Models
4.4 Evaluation Metrics
5 Results
@@ 6 Conclusion
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) [No. 2022R1A3B1077720], Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [NO. RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University), NO. RS-2022-II220959], the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2024, Samsung Electronics (IO221213-04119-01), and a grant from the Yang Young Foundation.
@@ Appendix A Appendix
A.1 Additional Details on EdiText Algorithms
In this section, we provide additional details on how the target attribute $c$ and the self-conditioning vector $\hat{x}$ are used in our algorithms, and our integrated editing methodology, building on the algorithms introduced in Sections 3.2 and 3.3. Our diffusion model, denoted $x_{\theta}$, takes as input not only the noisy sample $x_{t}$ at timestep $t$, but also (1) the target attribute $c$ and (2) a self-conditioning vector $\hat{x}$. Here, $c$ represents the label of the desired editing direction—such as toxifying or detoxifying in our toxicity task. Following the official LD4LG implementation, the model learns an embedding for $c$ during training on a toxicity dataset. This learned attribute embedding is then added with the timestep embedding to condition the diffusion model for class-conditional denoising, enabling generation or editing toward the specified attribute.
The self-conditioning vector $\hat{x}$ is treated according to the self-conditioning framework: it can be either a null vector or the prediction from a previous timestep, which the model uses to refine subsequent predictions. In our approach, consistent with LD4LG, this self-conditioning vector is handled internally as a separate context from the timestep or the target attribute embedding.
Algorithms 1 and 2 detail the coarse-level and fine-level editing procedures in EdiText, respectively. Given a reference text $w_{ref}$, we encode it using a language encoder E to obtain $x_{0}$. In the coarse-level editing (Algorithm 1), we apply a forward diffusion process on $x_{0}$ to obtain a perturbed reference $x_{t_{CE}}$, which then undergoes editing. In the fine-level editing (Algorithm 2), we utilize $x_{0}$ for $\hat{x}$ (the self-conditioning vector) as a condition during denoising, further refining the edits.
@@ A.3 Additional Results
A.4 Extension to Context-Conditional EdiText
A.5 Usage of AI Assistant